Aspirin (C9H8O4) is prepared by reacting salicylic acid (C7H6O3) with acetic anhydride (C4H6O3) according to the balanced equation:

C7H6O3(s) + C4H6O3(l) → C9H8O4(s) + C2H4O2(l)

A student is asked to prepare 90.08 g of aspirin. She is told that salicylic acid is the limiting reagent and that the reaction typically has a 53.9% yield. She is also told to use a 42.0% excess of acetic anhydride. How many grams of acetic anhydride should she use? Round your final answer to three significant figures.
1
Expert's answer
2019-09-30T05:21:49-0400

n(C9H8O4)=90.08g/180.157(g/mol)=0.5 moles

n(C4H6O3)=n(C9H8O4)=0.5 moles

0.5 moles*1.42=0.72 moles

m(C4H6O3)=0.72*102.1(g/mol)=73.5g